Hi, the combined land area of of the countries is 188,973 square kilometers. Country A is larger by 373 square kilometers. Determine the land area of each country. Thanks for any help, Tina
Answer by fractalier(6550)

Call them countries A and B.
A + B = 188973
and
A = B + 373
Substitute this into the first equation and get
(B + 373) + B = 188973
2B + 373 = 188973
Subtract 373 from both sides and get
2B = 188600
Now divide by two
B = 94300 sq km
A = B+373 = 94673 sq km